Luxed Construction Ltd (LCL) is a large service firm with two products in its product mix – SP General and SP Special. LCL’s 2018 books contain the following pieces of information (all budgeted data in the table below):

Activities                                 Total Costs          Applicable cost drivers

Clients Inquiries                         $201,000                     8,900 Inquiry hours

Clients billing                             $239,500                871,500 bills

Clients credit check                    $124,900                  63,000 accounts

Clients correspondence               $ 31,900                 13,500 letters

   Total costs                               $597,300

The two products’ consume LCL’s resources (i.e. use the business ‘activities’) in the following manner to run their respective day to day business affairs (all actual data in the table below):

                                                   SP General                  SP Special

Clients Inquiries                           2,900 hours                 6,000 hours

Clients billing                           271,500 bills                  600,000 bills

Clients credit check                   23,000 accounts             40,000 accounts

Clients correspondence                5,500 letters                    8,000 letters

Brand managers for SP General and SP Specials are paid annual retainer of $129,000 and $141,000, respectively.

Required: Work out the following showing all necessary calculations/working out (round the final answers to the nearest dollar; round individual indirect cost rates to 2 decimal places):

  1. The total cost to be assigned to SP General?
  2. The total cost to be assigned to SP Special?
  3. Briefly explain how the ABC system resulted in a more accurate costs assignment to each of the two products (respond in about 70 words)?
Answer #1

TOTAL COST ASSIGNED TO SP GENERAL :

Clients inquiries = 201000*2900/8900

= 65494.38

Clients billing = 239500*271500/871500

= 74611.88

Clients credit check = 124900*23000/63000

= 45598.41

Clients correspondence = 31900*5500/13500

= 12996.30

Brand Manager's salary = 129000

Therefore, TOTAL COST = 65494.38 + 74611.88 + 45598.41 + 12996.30 + 129000

= $ 327700.97

TOTAL COST ASSIGNED TO SP SPECIAL :

Clients inquiries = 201000*6000/8900

= 135505.62

Clients billing = 239500*600000/871500

= 164888.12

Clients credit check = 124900*40000/63000

= 79301.59

Clients correspondence = 31900*8000/13500

= 18903.70

Brand Manager's salary = 141000

Therefore, TOTAL COST = 135505.62 + 164888.12 + 79301.59 + 18903.70 + 141000

= $ 539599.03

Total cost assigned to SP GENERAL = $ 327700.97

Total cost assigned to SP SPECIAL = $ 539599.03